# 37N35, 65F55, 93A15, 93B15, 93C57

### v1

@@ 6 Numerical Results
We demonstrate the proposed $\mathsf{soQuadpvBT}$ method on three benchmark examples of second-order dynamical systems from the literature. The reported numerical experiments have been performed on a MacBook Air with 8 GB of RAM and an Apple M2 processor running macOS Ventura version 13.4 with MATLAB 23.2.0.2515942 (R2023b) Update 7. The source codes, data, and results of the numerical experiments are available at [35].
@@ 6.1 Experimental setup
is the intrusive position-velocity balanced truncation method for second-order systems from [34] and discussed in Section 2.3;
@@
The intrusive $\mathsf{sopvBT}$ is included as a point of comparison for the data-driven aspect of $\mathsf{soQuadpvBT}$. The intrusive and data-driven first-order methods $\mathsf{BT}$ and $\mathsf{foQuadBT}$ are included for the comparison between our proposed structured approach and unstructured models. The software package MORLAB version 6.0 [12, 10] is used for computing the intrusive reduced models by $\mathsf{sopvBT}$ and $\mathsf{BT}$.
For the data-based approaches, we use an exponential trapezoidal quadrature rule to implicitly approximate the relevant Gramians in eqs. 12 and 9. While $\mathsf{soLoewner}$ does not typically use interpolation points derived from quadrature rules, we fix the sampling points across all the data-based approaches for consistency. For simplicity, an equal number of left and right quadrature nodes is used so that $J=K=N$, and we assume that the nodes are symmetrically distributed according to eq. 30. In each example, the quadrature rules will be applied in chosen intervals along the imaginary axis.
@@ 6.2 Butterfly Gyroscope
[…]
@@ 6.2.1 Known damping coefficients
[…]
Surrogate models of order $r=10$ are computed using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For the data-based approaches, we use $N=200$ left and right quadrature nodes. These are chosen to be interwoven logarithmically spaced points in the interval $-\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and symmetrically distributed according to eq. 30. We use the realification process described in Section 5.1 for the $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ reduced models; for $\mathsf{foQuadBT}$, a similar procedure from [21, Sec. 4.1] is applied. The frequency response and pointwise relative errors eq. 35 of the reduced-order models are presented in Figure 1. We see that the data-driven approaches $\mathsf{soQuadpvBT}$ and $\mathsf{foQuadBT}$ are visibly indistinguishable from their intrusive counterparts $\mathsf{soBT}$ and $\mathsf{BT}$. Also, we observe that each of the approaches that preserve the underlying second-order structure, namely $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{sopvBT}$, provide very satisfactory approximations and are able to capture the response peaks of the full-order transfer function. By comparison, the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$ perform roughly two orders of magnitude worse throughout the entire frequency range, and entirely miss the response between $10^{5}$ and $10^{6}$ rad/s. This results from the fact that the first-order approximations having the same dimension as the second-order approximations cannot provide the same level of accuracy, regardless of whether that approximation is data-driven or intrusive. Table 1 shows the relative error measures eqs. 36a and 36b for the different methods. The structure-preserving methods clearly outperform the unstructured ones, and the $\mathsf{soQuadpvBT}$ and $\mathsf{soBT}$ reduced models produce very similar error values.
@@ 6.2.2 Estimating damping coefficients
Next, we employ the minimization routine from Section 5.2 to infer the Rayleigh damping coefficients from data, and use these estimated coefficients $\alpha_{\star}$ and $\beta_{\star}$ in the construction of the reduced-order damping matrix. For this set of experiments, we only consider $\mathsf{soQuadpvBT}$ to highlight the effects of the estimated coefficients on the reduced model error. The data $\boldsymbol{g}_{k}$ used to infer the damping coefficients are the evaluations of the transfer function $\boldsymbol{G}(\mathfrak{i}\vartheta_{k})$ and $\boldsymbol{G}(\mathfrak{i}\zeta_{k})$ at the left and right quadrature points. For solving the minimization problem eq. 31, we use the built-in MATLAB optimizer fminunc provided with the gradients eqs. 33a and 33b from Theorem 5. The internal minimization algorithm is the BFGS method with step and optimality tolerances both set to $10^{-8}$.
Figure 2: Frequency response and pointwise relative errors for the $r=10$ reduced-order $\mathsf{soQuadpvBT}$ models of the butterfly gyroscope using the true and estimated damping coefficients. The reduced-order models with inferred damping coefficients produce nearly identical results compared to the model with known parameters.
Table 2: Estimated Rayleigh damping coefficients $(\alpha_{\star},\beta_{\star})$ for the butterfly gyroscope example using different initializations $(\alpha_{i},\beta_{i})$.
$(\alpha_{1},\beta_{1})=(10^{-4},10^{-4})$ $(\alpha_{2},\beta_{2})=(0,0)$ $(\alpha_{3},\beta_{3})=(10^{-1},10^{-1})$
[…]
The minimization is applied to three different pairs of initial values for the Rayleigh damping coefficients, namely $(\alpha_{1},\beta_{1})=(10^{-4},10^{-4})$, $(\alpha_{2},\beta_{2})=(0,0)$ and $(\alpha_{3},\beta_{3})=(10^{-1},10^{-1})$, and the resulting local optima are shown in Table 2. Then, we used $\mathsf{soQuadpvBT}$ to compute four different reduced-order models based on the true damping parameters $(\alpha,\beta)=(0,10^{-6})$ and the three sets of inferred coefficients from Table 2. The frequency response and pointwise relative errors of these reduced models are all shown in Figure 2. The pointwise relative errors in Figure 2 reveal that the $\mathsf{soQuadpvBT}$ reduced models that use the estimated $\alpha_{\star}$ and $\beta_{\star}$ are nearly indistinguishable from the reduced model that incorporates the true $\alpha=0$ and $\beta=10^{-6}$. This shows that we can use $\mathsf{soQuadpvBT}$ to obtain suitable second-order reduced models as well as suitable damping coefficients using only given data.
@@ 6.2.3 Hermitian approach and stability preservation
Finally, we test the construction formulas from Theorem 2 to validate the stability result in Theorem 3. To make the system symmetric, we modify the input and position-output matrices to be $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}=\tfrac{1}{12}\boldsymbol{1}_{n}$, where $\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones of length $n$. For data generation, we use $N=200$ logarithmically spaced points $s_{i}$ in the interval $-\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and take the left and right quadrature nodes to be such that $\vartheta_{k}=-s_{k}$ and $\zeta_{j}=s_{j}$, which are symmetrically distributed according to eq. 30, and the weights satisfying $\eta_{i}=\varphi_{i}=\varrho_{i}$. Specifically, this organizational scheme implies that $s_{i}=-s_{i+1}$, for $i=1,3,\ldots,N-1$ and so the super and sub-diagonal entries of the matrices $\mathbb{M}$ and $\mathbb{K}$ require derivatives of $\boldsymbol{G}$ as in eq. 26.
Figure 3: Frequency response and pointwise relative errors for $r=20$ $\mathsf{soQuadpvBT}$ reduced-order models of the butterfly gyroscope using the data matrices from Theorem 1 and Theorem 2. The stability preserving model performs as well as the standard approach in terms of accuracy.
@@
For varying reduction orders $r=10,12,\ldots,20$, we compute reduced-order models via $\mathsf{soQuadpvBT}$ using the setup just described and the data matrices from Theorem 2, and for the comparison, we also compute reduced-order models using the setup of Section 6.2.1 and the data matrices in Theorem 1. In all cases, we employ the realification procedure from Section 5.1 to compute real-valued surrogates. The response and pointwise relative errors eq. 35 of the $r=20$ reduced-order are plotted in Figure 3. As illustrated by the figure, the $\mathsf{soQuadpvBT}$ reduced models using either set of data matrices provide very satisfactory approximations over the entire frequency range, and are nearly indistinguishable from each other. In Table 3, we report the stability properties of the reduced-order models. As expected from Theorem 3, the $\mathsf{soQuadpvBT}$ models constructed via Theorem 2 are all asymptotically stable by construction. Although not guaranteed, the $\mathsf{soQuadpvBT}$ reduced models constructed using Theorem 1 turned out to be asymptotically stable, as well.
@@ 6.3 Plate with tuned vibration absorbers
[…]
@@
$\mathsf{soQuadpvBT}$ $\mathsf{soLoewner}$ $\mathsf{foQuadBT}$
[…]
In this example, we only employ the data-driven methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{foQuadBT}$ due to the lack of available matrix equation solvers for the complex-valued coefficient matrices. Furthermore, we do not employ the realification process here because the underlying hysteretic damping leads to the full-order model being complex valued. For generating the data, we choose $N=250$ left and right quadrature nodes that are interwoven linearly spaced points in the interval $-2\pi\mathfrak{i}[0,250]\cup 2\pi\mathfrak{i}[0,250]$ and symmetrically distributed according to eq. 30. We use the aforementioned methods to compute reduced-order models of dimension $r=75$. The resulting frequency responses and pointwise relative errors are plotted in Figure 4. Similar to the previous example, the structure-preserving approaches $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ outperform the unstructured approach $\mathsf{foQuadBT}$ for most frequencies by about five orders of magnitude. The relative error measures reported in Table 4 also support this claim. We observe that all three reduced models struggle to produce a good approximation in the region around 48 Hz, which is due to the TVAs that heavily damp the frequency response about that region.
@@ 6.4 Mass-spring-damper network with velocity outputs
As the final example, we consider the mass-spring damper ($MSD$) chain from [43, Ex. 2]. This is a damped linear vibration system consisting of three rows of $d$ masses all connected to the right-hand side of a coupling mass $m_{0}$. The masses in the individual rows are connected by springs, and the coupling mass $m_{0}$ is connected to the fixed left-hand side of the base. The system order is $n=3d+1$. We construct a system with $d=300$ masses so that the full-order model dimension is $n=901$. We impose Rayleigh damping on the system for the parameters $f(s)=g(s)=0.002$, the input $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones, and we have the outputs $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$. We include this example to illustrate how our method performs when velocity outputs are incorporated.
Figure 5: Frequency response and pointwise relative errors for $r=20$ reduced-order models of the coupled $MSD$ network with a velocity output. The proposed $\mathsf{soQuadpvBT}$ approach yields similar performance to its intrusive counterpart $\mathsf{soBT}$.
Table 5: Relative $\mathcal{H}_{\infty}$ errors eq. 36a and $\mathcal{H}_{2}$ errors eq. 36b for $r=20$ reduced-order models of the $MSD$ chain. The smallest error is highlighted in boldface.
@@
[…]
We compute reduced-order models of size $r=20$ using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For generating the data, we choose $N=200$ left and right quadrature nodes as interwoven logarithmically spaced points in the interval $-\mathfrak{i}[10^{-3},10^{1}]\cup\mathfrak{i}[10^{-3},10^{1}]$, which are symmetrically distributed according to eq. 30. The frequency response of the full and reduced-order transfer functions as well as their pointwise relative errors eq. 35 are plotted in Figure 5. All methods produce reasonably good approximations to the full-order system response. As was the case for the previous two benchmark problems, we again observe that the structure-preserving methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, and $\mathsf{soBT}$ outperform the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$. This performance gap is particularly evident at the lower frequencies; all four methods exhibit roughly similar behavior from $1$ rad/s onward. The relative error measures are reported in Table 5, which shows that $\mathsf{sopvBT}$ performs the best overall.

### v2

@@ 6 Numerical Results
We demonstrate the proposed $\mathsf{soQuadpvBT}$ method on two benchmark examples of second-order dynamical systems from the literature. The reported numerical experiments have been performed on a MacBook Air with 8 GB of RAM and an Apple M2 processor running macOS Ventura version 13.4 with MATLAB 23.2.0.2515942 (R2023b) Update 7. The source codes, data, and results of the numerical experiments are available at [32].
@@ 6.1 Experimental setup
is the intrusive position-velocity balanced truncation for second-order systems from [31] and discussed in Section 2.3;
@@
The intrusive $\mathsf{sopvBT}$ is included as a point of comparison for the data-driven aspect of $\mathsf{soQuadpvBT}$. The intrusive and data-driven first-order methods $\mathsf{BT}$ and $\mathsf{foQuadBT}$ are included for the comparison between our proposed structured approach and unstructured models. The software package MORLAB version 6.0 [10] is used for computing the intrusive reduced models by $\mathsf{sopvBT}$ and $\mathsf{BT}$.
For the data-based approaches, we use an exponential trapezoidal quadrature rule to implicitly approximate the Gramians in eqs. 12 and 9. While $\mathsf{soLoewner}$ does not typically use interpolation points derived from quadrature rules, we fix the sampling points across all the data-based approaches for consistency. For simplicity, an equal number of left and right quadrature nodes is used so that $J=K=N$, and we assume that the nodes are symmetrically distributed according to eq. 30. In each example, the quadrature rules are applied in chosen intervals along the imaginary axis.
@@ 6.2 Butterfly Gyroscope
The first example that we consider is the butterfly gyroscope from the Oberwolfach Benchmark Collection [11], which models a vibrating mechanical gyroscope used for inertial navigation. The model itself is a second-order system eq. 1 with n=17 361n=17\,361 states, $m=1$ input and $p=12$ position outputs. The mass and stiffness matrices are symmetric, and the internal damping is described by Rayleigh damping with $f(s)=\alpha=0$ and $g(s)=\beta=10^{-6}$ according to eq. 5. The response behavior of interest occurs in the high-frequency ranges. We study this benchmark to illustrate how our proposed method applies to a system with Rayleigh damping eq. 6.
@@ 6.2.1 Known damping coefficients
$\operatorname{relerr}_{\mathcal{H}_{\infty}}$ $4.4215\texttt{e-}4$ $4.5263\texttt{e-}4$ $1.0498\texttt{e-}2$ $\boldsymbol{4.4202\texttt{e-}4}$ $1.0505\texttt{e-}2$
$\operatorname{relerr}_{\mathcal{H}_{2}}$ $8.0512\texttt{e-}4$ $8.3923\texttt{e-}4$ $2.5685\texttt{e-}2$ $\boldsymbol{8.0166\texttt{e-}4}$ $2.5297\texttt{e-}2$
@@
Surrogate models of order $r=10$ are computed using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For the data-based approaches, we use $N=200$ left and right quadrature nodes. These are chosen to be interwoven logarithmically spaced points in the interval $\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and symmetrically distributed according to eq. 30. We use the realification process described in Section 5.1 for the $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ reduced models; for $\mathsf{foQuadBT}$, a similar procedure from [19, Sec. 4.1] is applied. The frequency response and pointwise relative errors eq. 35 of the reduced-order models are presented in Figure 1. We see that the data-driven approaches $\mathsf{soQuadpvBT}$ and $\mathsf{foQuadBT}$ are visibly indistinguishable from their intrusive counterparts $\mathsf{soBT}$ and $\mathsf{BT}$. Also, we observe that each of the approaches that preserve the underlying second-order structure, namely $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{sopvBT}$, provide very satisfactory approximations and are able to capture the response peaks of the full-order transfer function. By comparison, the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$ perform roughly two orders of magnitude worse throughout the entire frequency range, and entirely miss the response between $10^{5}$ and $10^{6}$ rad/s. This results from the fact that the first-order approximations having the same dimension as the second-order approximations cannot provide the same level of accuracy, regardless of whether that approximation is data-driven or intrusive. Table 1 shows the relative error measures eqs. 36a and 36b for the different methods. The structure-preserving methods clearly outperform the unstructured ones, and the $\mathsf{soQuadpvBT}$ and $\mathsf{soBT}$ reduced models produce very similar error values.
@@ 6.2.2 Estimating damping coefficients
Next, we employ the minimization procedure from Section 5.2 to infer the Rayleigh damping coefficients from data, and use these estimated coefficients $\alpha_{\star}$ and $\beta_{\star}$ in the construction of the reduced-order models. For this set of experiments, we only consider $\mathsf{soQuadpvBT}$ to highlight the effects of the estimated coefficients on the reduced model error. The data $\boldsymbol{g}_{k}$ used to infer the damping coefficients are the evaluations of the transfer function $\boldsymbol{G}(\mathfrak{i}\vartheta_{k})$ and $\boldsymbol{G}(\mathfrak{i}\zeta_{k})$ at the left and right quadrature points. For solving the minimization problem eq. 31, we use the optimizer particleswarm provided by MATLAB’s global optimization toolbox. Once suitable minima $\alpha_{\star}$ and $\beta_{\star}$ are determined and a reduced model is computed by Algorithm 2 using these parameters, we employ the post-processing step described in Section 5.2 using the gradients eq. 33 from Theorem 5 to form the reduced-order damping term. For this post-processing step, the built-in MATLAB optimizer fminunc is used; the internal minimization algorithm is the BFGS method with step and optimality tolerances both set to $10^{-10}$.
Figure 2: Frequency response and pointwise relative errors for the $r=10$ reduced-order $\mathsf{soQuadpvBT}$ models of the butterfly gyroscope using the true and estimated damping coefficients.
$\mathcal{I}_{\alpha}^{1}\times\mathcal{I}_{\beta}^{1}=[0,10^{-8}]\times[10^{-7},10^{-5}]$ $\mathcal{I}_{\alpha}^{2}\times\mathcal{I}_{\beta}^{2}=[0,10^{-8}]\times[10^{-8},10^{-4}]$
Estimated $\alpha_{\star}$ $8.4361\texttt{e-}9$ $1.0000\texttt{e-}8$
Estimated $\beta_{\star}$ $9.8017\texttt{e-}9$ $3.8284\texttt{e-}5$
Table 2: Estimated Rayleigh damping coefficients $(\alpha_{\star},\beta_{\star})$ for the butterfly gyroscope example over different intervals $\mathcal{I}_{\alpha}^{i}$ and $\mathcal{I}_{\beta}^{i}$.
The minimization is applied over two different pairs of intervals, namely $\mathcal{I}_{\alpha}^{1}\times\mathcal{I}_{\beta}^{1}=[0,10^{-8}]\times[10^{-7},10^{-5}]$ and $\mathcal{I}_{\alpha}^{2}\times\mathcal{I}_{\beta}^{2}=[0,10^{-8}]\times[10^{-8},10^{-4}]$, and the resulting local optima are shown in Table 2. Three different $\mathsf{soQuadpvBT}$ reduced-order models are computed based on the true damping parameters $(\alpha,\beta)=(0,10^{-6})$ and the two sets of inferred coefficients from Table 2. The frequency response and pointwise relative errors of these reduced models are all shown in Figure 2. The pointwise relative errors in Figure 2 reveal that the $\mathsf{soQuadpvBT}$ reduced model that uses the estimated $\alpha_{\star}$ and $\beta_{\star}$ corresponding to $\mathcal{I}_{\alpha}^{1}$ and $\mathcal{I}_{\beta}^{1}$ are nearly indistinguishable from the reduced model that incorporates the true $\alpha=0$ and $\beta=10^{-6}$. On the other hand, the reduced model that uses the coefficients estimated by optimizing over $\mathcal{I}_{\alpha}^{2}$ and $\mathcal{I}_{\beta}^{2}$ performs a few orders of magnitude worse. This is because the particles get stuck in a stationary point of the objective function eq. 31. This shows that we can use $\mathsf{soQuadpvBT}$ to obtain suitable second-order reduced models as well as suitable damping coefficients using only the given data if we have a sufficiently good initial search interval.
@@ 6.2.3 Hermitian approach and stability preservation
Finally, we test the construction formulas from Theorem 2 to validate the stability result in Theorem 3. To make the system symmetric, we modify the input and position-output matrices to be $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}=\tfrac{1}{12}\boldsymbol{1}_{n}$, where $\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones of length $n$. For data generation, we use $N=200$ logarithmically spaced points $s_{i}$ in the interval $\mathfrak{i}[10^{4},10^{6}]\cup\mathfrak{i}[10^{4},10^{6}]$ and take the left and right quadrature nodes to be such that $\vartheta_{k}=-s_{k}$ and $\zeta_{j}=s_{j}$, which are symmetrically distributed according to eq. 30, and the weights satisfying $\xi_{i}=\varphi_{i}=\varrho_{i}$. Specifically, this organizational scheme implies that $s_{i}=-s_{i+1}$, for $i=1,3,\ldots,N-1$ and so the super and sub-diagonal entries of the matrices $\mathbb{M}$ and $\mathbb{K}$ require derivatives of $\boldsymbol{G}$ as in eq. 26.
Figure 3: Frequency response and pointwise relative errors for $r=20$ $\mathsf{soQuadpvBT}$ reduced-order models of the butterfly gyroscope using the data matrices from Theorem 1 and Theorem 2. The stability-preserving model performs as well as the standard approach in terms of accuracy.
@@
For varying reduction orders $r=10,12,\ldots,20$, we compute reduced-order models via $\mathsf{soQuadpvBT}$ using the setup just described and the data matrices from Theorem 2, and for the comparison, we also compute reduced-order models using the setup of Section 6.2.1 and the data matrices in Theorem 1. In all cases, we employ the realification procedure from Section 5.1 to compute real-valued surrogates. The response and pointwise relative errors eq. 35 of the $r=20$ reduced-order models are plotted in Figure 3. As illustrated by the figure, the $\mathsf{soQuadpvBT}$ reduced models using either set of data matrices provide very satisfactory approximations over the entire frequency range, and are nearly indistinguishable from each other. In Table 3, we report the stability properties of the reduced-order models. As expected from Theorem 3, the $\mathsf{soQuadpvBT}$ models constructed via Theorem 2 are all asymptotically stable by construction. Although not guaranteed, the $\mathsf{soQuadpvBT}$ reduced models constructed using Theorem 1 turned out to be asymptotically stable, as well.
@@ 6.3 Plate with tuned vibration absorbers
As a second example, we consider the model of the vibrational response of a strutted plate from [3, Sec. 4.2]. We refer the reader to [3, Sec. 4.2] for the material parameters and model details. The model is a second-order system eq. 1 with n=209 100n=209\,100 states, $m=1$ input and $p=1$ position output, and it uses the structural (hysteretic) damping formulation with $g(s)=(\mathfrak{i}\eta)/s$ for $\eta=10^{-3}$. The tuned vibration absorbers (TVAs) connected to the plate reduce the vibrational response in the frequency region about $48$ Hz.
In this example, we only employ the data-driven methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$ and $\mathsf{foQuadBT}$ and omit the intrusive model reduction counterparts due to the lack of available solvers for the algebraic Lyapunov equations with complex-valued coefficient matrices. Furthermore, we do not employ the realification process here because the underlying hysteretic damping leads to the full-order model being complex valued. For generating the data, we choose $N=250$ left and right quadrature nodes that are interwoven linearly spaced points in the interval $2\pi\mathfrak{i}[0,250]\cup 2\pi\mathfrak{i}[0,250]$ and symmetrically distributed according to eq. 30. We use the aforementioned methods to compute reduced-order models of dimension $r=75$. The resulting frequency responses and pointwise relative errors are plotted in Figure 4. Similar to the previous example, the structure-preserving approaches $\mathsf{soQuadpvBT}$ and $\mathsf{soLoewner}$ outperform the unstructured approach $\mathsf{foQuadBT}$ for most frequencies by about five orders of magnitude. The relative error measures reported in Table 4 also support this claim. We observe that all three reduced models struggle to produce a good approximation in the region around $48$ Hz, which is due to the TVAs that heavily damp the frequency response about that region.
@@
$\mathsf{soQuadpvBT}$ $\mathsf{soLoewner}$ $\mathsf{foQuadBT}$
$\operatorname{relerr}_{\mathcal{H}_{\infty}}$ $2.0148\texttt{e-}9$ $\boldsymbol{1.8776\texttt{e-}9}$ $8.7564\texttt{e-}8$
$\operatorname{relerr}_{\mathcal{H}_{2}}$ $2.7231\texttt{e-}10$ $\boldsymbol{2.5159\texttt{e-}10}$ $1.3176\texttt{e-}8$
@@ 6.4 Mass-spring-damper network with velocity outputs
As the final example, we consider the mass-spring damper ($MSD$) chain from [40, Ex. 2]. This is a damped linear vibration system consisting of three rows of $d$ masses all connected to the right-hand side of a coupling mass $m_{0}$. The masses in the individual rows are connected by springs, and the coupling mass $m_{0}$ is connected to the fixed left-hand side of the base. The system order is $n=3d+1$. We construct a system with $d=300$ masses so that the full-order model dimension is $n=901$. We impose Rayleigh damping on the system for the parameters $f(s)=g(s)=0.002$, the input $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones, and we have the outputs $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{1}_{n}\in\mathbb{R}^{n}$. We include this example to illustrate how our method performs when velocity outputs are incorporated.
Figure 5: Frequency response and pointwise relative errors for $r=20$ reduced-order models of the coupled MSDMSD network with a velocity output. The proposed $\mathsf{soQuadpvBT}$ approach yields similar performance to its intrusive counterpart $\mathsf{soBT}$.
@@
$\operatorname{relerr}_{\mathcal{H}_{\infty}}$ $1.2550\texttt{e-}3$ $2.9718\texttt{e-}3$ $4.2038\texttt{e-}3$ $\boldsymbol{9.3830\texttt{e-}4}$ $3.6199\texttt{e-}3$
$\operatorname{relerr}_{\mathcal{H}_{2}}$ $1.0782\texttt{e-}3$ $1.5801\texttt{e-}3$ $4.2038\texttt{e-}3$ $\boldsymbol{8.5111\texttt{e-}4}$ $4.8201\texttt{e-}3$
Table 5: Relative $\mathcal{H}_{\infty}$ errors eq. 36a and $\mathcal{H}_{2}$ errors eq. 36b for $r=20$ reduced-order models of the MSDMSD chain. The smallest error is highlighted in boldface.
We compute reduced-order models of size $r=20$ using $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, $\mathsf{foQuadBT}$, $\mathsf{sopvBT}$ and $\mathsf{BT}$. For generating the data, we choose $N=200$ left and right quadrature nodes as interwoven logarithmically spaced points in the interval $\mathfrak{i}[10^{-3},10^{1}]\cup\mathfrak{i}[10^{-3},10^{1}]$, which are symmetrically distributed according to eq. 30. The frequency response of the full and reduced-order transfer functions as well as their pointwise relative errors eq. 35 are plotted in Figure 5. All methods produce reasonably good approximations to the full-order system response. As was the case for the previous two benchmark problems, we again observe that the structure-preserving methods $\mathsf{soQuadpvBT}$, $\mathsf{soLoewner}$, and $\mathsf{soBT}$ outperform the unstructured methods $\mathsf{foQuadBT}$ and $\mathsf{BT}$. This performance gap is particularly evident at the lower frequencies; all four methods exhibit roughly similar behavior from $1$ rad/s onward. The relative error measures are reported in Table 5, which shows that $\mathsf{sopvBT}$ performs the best overall.